# Close-in compact super-Earth systems emerging from resonant chains: slow destabilization by unseen remnants of formation

### v1

@@ 4.4 Is something missing?
Interestingly, the strength of the observed Kepler dichotomy is not reached even in our highest mass simulations. In other words, violent instabilities that produce enough single transiting planets also tend to make too widely-spaced systems (e.g., Fig 10) and vice versa (Fig 9). Our 40straight suite was a good match to both the observed multiplicity and period ratio distribution (Fig 13) but predominately makes planets of mass (Fig 7), seemingly too high compared to the typical Kepler super-Earth or sub-Neptune of (He2020; Rogers2021). This problem also appears in Izidoro2021a, who tested multiple pebble accretion models with different mass flux. Their sole model that matches the observed multiplicity distribution (Model-I, in their notation), has a typical planet mass of .
@@
One potential way to bridge this gap is an additional source of inclination excitation. Spalding2016 showed that mutual inclinations can be excited in multiplanet systems through the quadrupolar precession of a primordially misaligned host star. During spindown, a secular resonance can be excited that tilts the outer planet to a non-transiting configuration. Although this process can itself trigger a dynamical instability (Spalding2018), the quiescent version appears a promising mechanism to raise mutual inclinations in some systems without requiring large masses or orbital spacings between planets.
@@
We also note that the very steep decline in the observed incidence of resonance between the first and second bin of Fig 14 is not well reproduced in our simulations. While Dai2024 presented a compelling argument for a population trend, there is reason to believe that the quantitative details are not set in stone. For one, Dai2024 did not attempt to debias the sample. Young transiting planets present numerous difficulties in detection and confirmation due to stellar activity and inhomogeneous distribution on the sky. Furthermore, near-resonant planets experience transit timing variations that smear transits after phase-folding and impede detection (Leleu2021a). Neither of these effects nor their combination has been carefully treated in the literature. We expect that the PLATO mission, which is intended to provide a robust statistical sample as well as accurately measured ages for many planet hosts, will greatly refine the observational constraints and potentially clarify whether there is in fact a preferred time of instability.
@@ 5 Conclusion
Pressure bumps and their resulting migration traps are effective ways to increase the rate of instabilities compared to purely power-law disks. There is frequently an outer resonant chain that is unstable on long timescales and passes its instability to the inner chain, which would have been stable on its own. These outer companions are not easily observable with current transiting or radial velocity techniques, but may be detectable by microlensing. The role of eccentric outer planets in triggering instabilities in an inner chain has also been studied simultaneously and independently by Ogihara2025.
Compact planetary systems continuously evolve over hundreds of Myr, or billions of orbits. Comparing planet formation models that end at 100 Myr or less to mature systems with ages 1–10 Gyr can be deeply misleading. The common issue that many population synthesis models produce too many resonant systems may simply be an artifact of ending N-body integrations too early. If the high fraction of mean-motion resonance in young systems reported by Dai2024 is confirmed by further observations, this may be a correct prediction of these models.
@@
As a whole, our findings strongly support the “breaking-the-chains” model as the dominant pathway for the formation of systems of small planets within 100 d. It has become increasingly clear that the architectures of planetary systems are not set at the moment of disk dissipation (Berger2020; Yang2023). Current and upcoming observations with TESS and PLATO, especially those of young systems and with well-measured host star ages, will help us map the evolution of planetary systems over their lifetimes. The onus then lies on theorists to produce models of dynamical evolution extending beyond the first 1–10% of the lifetimes of typical systems that can be used to reliably infer details of their initial conditions and formation. This will be a formidable challenge for traditional N-body methods, but continuing advancements in accurate semi-analytical models may be a promising path forward (Kimura2025; Kokubo2025).
@@ Acknowledgements.
We thank Alessandro Morbidelli, Masahiro Ogihara, Masanobu Kunitomo, and Konstantin Batygin for insightful discussions. This work was funded by the ERC project N. 101019380 “HolyEarth.”Appendix A Numerical tests

### v2

@@ 4.4 Is something missing?
Interestingly, the strength of the observed Kepler dichotomy is not reached even in our highest mass simulations. In other words, violent instabilities that produce enough single transiting planets also tend to make too widely-spaced systems (e.g., Fig 6) and vice versa (Fig 5). Our 40straight suite was a good match to both the observed multiplicity and period ratio distribution (Fig 9) but predominately makes planets of mass (Fig 17), seemingly too high compared to the typical Kepler super-Earth or sub-Neptune of (He et al., 2020; Rogers and Owen, 2021). This problem also appears in Izidoro et al. (2021), who tested multiple pebble accretion models with different mass flux. Their sole model that matches the observed multiplicity distribution (Model-I, in their notation), has a typical planet mass of .
@@
One potential way to bridge this gap is an additional source of inclination excitation. Spalding and Batygin (2016) showed that mutual inclinations can be excited in multiplanet systems through the quadrupolar precession of a primordially misaligned host star. During spindown, a secular resonance can be excited that tilts the outer planet to a non-transiting configuration. Although this process can itself trigger a dynamical instability (Spalding et al., 2018), the quiescent version appears a promising mechanism to raise mutual inclinations in some systems without requiring large masses or orbital spacings between planets.
Orbital eccentricities are more difficult to compare to models because they are not directly measurable from transit observations. Transit timing variations provide more information on the eccentricity through the mean-motion resonant dynamics. Yee et al. (2021) showed that TTV systems have eccentricities several times lower than that required for stability, suggesting some amount of dissipation that damped eccentricity following a giant impact phase. However, TTV systems represent a biased sample of near-resonant planets, which, in the context of the breaking-the-chains model, are less likely to have undergone an instability that excites eccentricity (Goldberg and Batygin, 2022).
Transit duration measurements, on the other hand, can constrain the average eccentricity of an entire population of planets and have indicated that Kepler multiplanet systems have typical eccentricities of (Van Eylen et al., 2019; Mills et al., 2019). Gilbert et al. (2025) refined this strategy by dividing the population in planet radius bins and found that small planets () have mean eccentricity of 0.06 in single transiting systems and 0.03 in multitransiting systems. After applying the synthetic observational bias, we find mean eccentricities of 0.06 and 0.05 in single and multitransiting systems in the 20flat simulations and 0.09 and 0.07 in the 40steep simulations. This is consistent with a small degree of eccentricity damping over the lifetime of the system after a dynamical instability, such as that provided by tidal dissipation (Delisle and Laskar, 2014).
This model predicts that the youngest observed systems should be predominantly chains of planetary deep in resonance. Although Dai et al. (2024) identified a surplus of systems near resonance according to an orbital period based criterion, most of these systems do not have sufficient dynamical constraints to determine whether they are actually resonant according to a dynamical definition that requires libration of the resonant angle within a separatrix-bounded region. Hu et al. (2025) studied several of these systems in detail and found several planet pairs that, while classified as near-resonant in terms of period ratio, are sufficiently distant from exact commensurability that their resonant angles are likely to be circulating. As near-resonance in our synthetic systems typically implies resonant libration, these intriguing observational findings imply additional physics not captured in our simulations during the disk phase or shortly after disk dissipation that can modify the resonant state.
@@
Finally, we note that the very steep decline in the observed incidence of resonance between the first and second bin of Fig 10 is not well reproduced in our simulations. While Dai et al. (2024) presented a compelling argument for a population trend, there is reason to believe that the quantitative details are not set in stone. For one, Dai et al. (2024) did not attempt to debias the sample. Young transiting planets present numerous difficulties in detection and confirmation due to stellar activity and inhomogeneous distribution on the sky. Furthermore, near-resonant planets experience transit timing variations that smear transits after phase-folding and impede detection (Leleu et al., 2021b). Neither of these effects nor their combination has been carefully treated in the literature. We expect that the PLATO mission, which is intended to provide a robust statistical sample as well as accurately measured ages for many planet hosts, will greatly refine the observational constraints and potentially clarify whether there is in fact a preferred time of instability.
@@ 5 Conclusion
Pressure bumps and their resulting migration traps are effective ways to increase the rate of instabilities compared to purely power-law disks. There is frequently an outer resonant chain that is unstable on long timescales and passes its instability to the inner chain, which would have been stable on its own. These outer companions are not easily observable with current transiting or radial velocity techniques, but may be detectable by microlensing. The role of eccentric outer planets in triggering instabilities in an inner chain has also been studied simultaneously and independently by Ogihara and Kunitomo (2026).
Compact planetary systems continuously evolve over hundreds of Myr, or billions of orbits. Comparing planet formation models that end at 100 Myr or less to mature systems with ages 1–10 Gyr can be deeply misleading. The common issue that many population synthesis models produce too many resonant systems may simply be an artifact of ending N-body integrations too early. If the high fraction of mean-motion resonance in young systems reported by Dai et al. (2024) is confirmed by further observations, this may be a correct prediction of these models.
@@
As a whole, our findings strongly support the “breaking-the-chains” model as the dominant pathway for the formation of systems of small planets within 100 d. It has become increasingly clear that the architectures of planetary systems are not set at the moment of disk dissipation (Berger et al., 2020; Yang et al., 2023). Current and upcoming observations with TESS and PLATO, especially those of young systems and with well-measured host star ages, will help us map the evolution of planetary systems over their lifetimes. The onus then lies on theorists to produce models of dynamical evolution extending beyond the first 1–10% of the lifetimes of typical systems that can be used to reliably infer details of their initial conditions and formation. This will be a formidable challenge for traditional N-body methods, but continuing advancements in accurate semi-analytical models may be a promising path forward (Kimura et al., 2025; Kokubo et al., 2025).
@@ Acknowledgements.
We are grateful to referee Dan Tamayo for useful comments that improved the paper. We thank Alessandro Morbidelli, Masahiro Ogihara, Masanobu Kunitomo, and Konstantin Batygin for insightful discussions. This work was funded by the ERC project N. 101019380 “HolyEarth.”. A. P. is supported by the French government, through the Investments in the Future project managed by the National Research Agency (ANR) with the reference number ANR-15-IDEX-01.References
- Pushing planets into an inner cavity by a resonant chain. A&A 648, pp. A69. External Links: ISSN 0004-6361, Document Cited by: §2.
- An Analytic Criterion for Turbulent Disruption of Planetary Resonances. AJ 153, pp. 120. External Links: ISSN 0004-6256, Document Cited by: §4.3.
- Formation of rocky super-earths from a narrow ring of planetesimals. Nature Astronomy. External Links: ISSN 2397-3366, Document Cited by: §1, §2, §3.1.
- The Gaia-Kepler Stellar Properties Catalog. II. Planet Radius Demographics as a Function of Stellar Mass and Age. AJ 160, pp. 108. External Links: ISSN 0004-6256, Document Cited by: §5.
- Atmospheric mass-loss due to giant impacts: the importance of the thermal component for hydrogen-helium envelopes. MNRAS 485, pp. 4454–4463. External Links: ISSN 0035-8711, Document Cited by: §4.3.
- Diversity of disc viscosities can explain the period ratios of resonant and non-resonant systems of hot super-Earths and mini-Neptunes. A&A 692, pp. A246. External Links: ISSN 0004-6361, Document Cited by: §4.2.
- Efficient Geometric Probabilities of Multi-Transiting Exoplanetary Systems from CORBITS. ApJ 821, pp. 47. External Links: ISSN 0004-637X, Document Cited by: §3.4.
- Young Planets around Young Accreting Stars: I. Migration and Inner Stalling Orbits. arXiv e-prints, pp. arXiv:2505.03701. External Links: Document, 2505.03701, ADS entry Cited by: §4.3.
- A hybrid symplectic integrator that permits close encounters between massive bodies. MNRAS 304, pp. 793–799. External Links: ISSN 0035-8711, Document Cited by: §2.
- Tidal interactions shape period ratios in planetary systems with three-body resonant chains. A&A 677, pp. A160. External Links: ISSN 0004-6361, Document Cited by: §4.2.
- Capture and escape of planetary mean-motion resonances in turbulent discs. MNRAS 540, pp. 1998–2007. External Links: ISSN 0035-8711, Document Cited by: §4.3.
- On the evolution of multiple protoplanets embedded in a protostellar disc. A&A 450, pp. 833–853. External Links: ISSN 0004-6361, Document Cited by: §1, §2.
- Three-dimensional simulations of multiple protoplanets embedded in a protostellar disc. A&A 482, pp. 677–690. External Links: ISSN 0004-6361, Document Cited by: §1, §2.
- The Prevalence of Resonance Among Young, Close-in Planets. AJ 168, pp. 239. External Links: ISSN 0004-6256, Document Cited by: §1, Figure 10, §3.3, §3.4, §4.4, §4.4, 3rd item.
- TOI-1136 is a Young, Coplanar, Aligned Planetary System in a Pristine Resonant Chain. AJ 165, pp. 33. External Links: ISSN 0004-6256, Document Cited by: §4.2.
- First-order Resonance Overlap and the Stability of Close Two-planet Systems. ApJ 774, pp. 129. External Links: ISSN 0004-637X, Document Cited by: §4.1.
- Tidal dissipation and the formation of Kepler near-resonant planets. A&A 570, pp. L7. External Links: ISSN 0004-6361, Document Cited by: §4.4.
- Atmosphere loss in planet-planet collisions. MNRAS 496, pp. 1166–1181. External Links: ISSN 0035-8711, Document Cited by: §4.3.
- The New Generation Planetary Population Synthesis (NGPPS). I. Bern global model of planet formation and evolution, model tests, and emerging planetary systems. A&A 656, pp. A69. External Links: ISSN 0004-6361, Document Cited by: §1.
- The New Generation Planetary Population Synthesis (NGPPS). II. Planetary population of solar-like stars and overview of statistical results. A&A 656, pp. A70. External Links: ISSN 0004-6361, Document Cited by: §3.4.
- The origins of nearly coplanar, non-resonant systems of close-in super-Earths. MNRAS 497, pp. 2493–2500. External Links: Document Cited by: §3.2.
- 3D Radiation Nonideal Magnetohydrodynamical Simulations of the Inner Rim in Protoplanetary Disks. ApJ 835, pp. 230. External Links: ISSN 0004-637X, Document Cited by: §2.
- The False Positive Rate of Kepler and the Occurrence of Planets. ApJ 766, pp. 81. External Links: ISSN 0004-637X, Document Cited by: §1.
- Planets larger than Neptune have elevated eccentricities. PNAS 122, pp. e2405295122. External Links: ISSN 0027-8424, Document Cited by: §4.4.
- Seven temperate terrestrial planets around the nearby ultracool dwarf star TRAPPIST-1. Nature 542, pp. 456–460. External Links: ISSN 0028-0836, Document Cited by: §3.
- A criterion for the stability of planets in chains of resonances. Icarus 388, pp. 115206. External Links: ISSN 0019-1035, Document Cited by: §3.2, §4.1.
- A Tidal Origin for a Three-body Resonance in Kepler-221. AJ 162, pp. 16. External Links: ISSN 0004-6256, Document Cited by: §4.2, §4.2.
- Architectures of Compact Super-Earth Systems Shaped by Instabilities. AJ 163, pp. 201. External Links: ISSN 0004-6256, Document Cited by: §1, §1, §3.4, §4.4.
- Overstable Librations can Account for the Paucity of Mean Motion Resonances among Exoplanet Pairs. AJ 147, pp. 32. External Links: ISSN 0004-6256, Document Cited by: §3.1.
- Kepler-discovered Multiple-planet Systems near Period Ratios Suggestive of Mean-motion Resonances Are Young. AJ 167, pp. 55. External Links: ISSN 0004-6256, Document Cited by: §1.
- Widespread disruption of resonant chains during protoplanetary disk dispersal. The Open Journal of Astrophysics 7, pp. 61. External Links: ISSN 2565-6120, Document Cited by: §1.
- Architectures of Exoplanetary Systems. III. Eccentricity and Mutual Inclination Distributions of AMD-stable Planetary Systems. AJ 160, pp. 276. External Links: ISSN 0004-6256, Document Cited by: §1, §4.4.
- Switching integrators reversibly in the astrophysical N-body problem. MNRAS 522, pp. 4639–4648. External Links: ISSN 0035-8711, Document Cited by: §2.
- Unexpected Near-Resonant and Metastable States of Young Multiplanet Systems. ApJ 995, pp. 206. External Links: ISSN 0004-637X, Document Cited by: §1, §4.4.
- Fundamental limits from chaos on instability time predictions in compact planetary systems. MNRAS 491, pp. 5258–5267. External Links: ISSN 0035-8711, Document Cited by: Appendix B.
- Formation of planetary systems by pebble accretion and migration. Hot super-Earth systems from breaking compact resonant chains. A&A 650, pp. A152. External Links: ISSN 0004-6361, Document Cited by: §1, §1, §3.4, §4.4.
- Breaking the chains: hot super-Earth systems from migration and disruption of compact resonant chains. MNRAS 470 (2), pp. 1750–1770. External Links: ISSN 0035-8711, Document Cited by: §1, §1, §2, §3.4, §3.4, §3, §4.3.
- The Exoplanet Radius Valley from Gas-driven Planet Migration and Breaking of Resonant Chains. ApJ 939, pp. L19. External Links: ISSN 0004-637X, Document Cited by: §4.3.
- Can Planetary Instability Explain the Kepler Dichotomy?. ApJ 758, pp. 39. External Links: ISSN 0004-637X, Document Cited by: §3.4.
- Mean motion resonance capture in the context of type I migration. A&A 669, pp. A44. External Links: ISSN 0004-6361, Document Cited by: §3.2, §4.2.
- Semianalytical Model for the Dynamical Evolution of Planetary Systems. II. Application to Systems Formed by a Planet Formation Model. ApJ 989, pp. 117. External Links: ISSN 0004-637X, Document Cited by: §4.3, §5.
- A Scaling Law for the Orbital Architecture of Planetary Systems Formed by Gravitational Scattering and Collisions. ApJ 991, pp. L49. External Links: ISSN 0004-637X, Document Cited by: §5.
- Dispersal of protoplanetary discs by the combination of magnetically driven and photoevaporative winds. MNRAS 492, pp. 3849–3858. External Links: ISSN 0035-8711, Document Cited by: §2.
- Intra-system uniformity: a natural outcome of dynamical sculpting. MNRAS 525, pp. L66–L71. External Links: ISSN 0035-8711, Document Cited by: §1.
- High order symplectic integrators for perturbed Hamiltonian systems. CeMDA 80, pp. 39–62. External Links: ISSN 0923-2958, Document Cited by: Appendix B.
- Six transiting planets and a chain of Laplace resonances in TOI-178. A&A 649, pp. A26. External Links: ISSN 0004-6361, Document Cited by: §3.
- Alleviating the transit timing variation bias in transit surveys. I. RIVERS: Method and detection of a pair of resonant super-Earths around Kepler-1705. A&A 655, pp. A66. External Links: ISSN 0004-6361, Document Cited by: §4.4.
- Resonant sub-Neptunes are puffier. A&A 687, pp. L1. External Links: ISSN 0004-6361, Document Cited by: §1.
- Dynamical rearrangement of super-Earths during disk dispersal. I. Outline of the magnetospheric rebound model. A&A 601, pp. A15. External Links: ISSN 0004-6361, Document Cited by: §1.
- TRACE: a code for time-reversible astrophysical close encounters. MNRAS 533, pp. 3708–3723. External Links: ISSN 0035-8711, Document Cited by: §2.
- Disk Surface Density Transitions as Protoplanet Traps. ApJ 642, pp. 478–487. External Links: ISSN 0004-637X, Document Cited by: §2.
- Evidence for a Nondichotomous Solution to the Kepler Dichotomy: Mutual Inclinations of Kepler Planetary Systems from Transit Duration Variations. AJ 162, pp. 166. External Links: ISSN 0004-6256, Document Cited by: §3.4.
- The California-Kepler Survey. VIII. Eccentricities of Kepler Planets and Tentative Evidence of a High-metallicity Preference for Small Eccentric Planets. AJ 157, pp. 198. External Links: ISSN 0004-6256, Document Cited by: §4.4.
- The Exoplanet Population Observation Simulator. I. The Inner Edges of Planetary Systems. AJ 156, pp. 24. External Links: ISSN 0004-6256, Document Cited by: §2.
- Numerical optimization. 2nd ed edition, Springer Series in Operations Research, Springer, New York. External Links: ISBN 978-0-387-30303-1, LCCN QA402.5 .N62 2006 Cited by: §3.3.
- Formation and Disruption of Resonant Chains of Super-Earths: Secular Perturbations from Outer Eccentric Embryos. ApJ 996, pp. 91. External Links: ISSN 0004-637X, Document Cited by: §4.1, 2nd item.
- Super-Earth Formation with Slow Migration from a Ring in an Evolving Peaked Disk Compatible with Terrestrial Planet Formation. ApJ 972, pp. 181. External Links: ISSN 0004-637X, Document Cited by: §1, §4.3.
- Mapping out the parameter space for photoevaporation and core-powered mass-loss. MNRAS 528, pp. 1615–1629. External Links: ISSN 0035-8711, Document Cited by: §1.
- A torque formula for non-isothermal type I planetary migration - I. Unsaturated horseshoe drag. MNRAS 401, pp. 1950–1964. External Links: ISSN 0035-8711, Document Cited by: §2.
- A torque formula for non-isothermal Type I planetary migration - II. Effects of diffusion. MNRAS 410, pp. 293–303. External Links: ISSN 0035-8711, Document Cited by: §2.
- The Role of Magnetospheric Rebound in Breaking Resonant Chains of Super-Earths and Mini-Neptunes. ApJ 994, pp. 33. External Links: ISSN 0004-637X, Document Cited by: §1.
- Three body resonances in close orbiting planetary systems: tidal dissipation and orbital evolution. International Journal of Astrobiology 14, pp. 291–304. External Links: Document Cited by: §4.2.
- Prevalence of Earth-size planets orbiting Sun-like stars. PNAS 110, pp. 19273–19278. External Links: ISSN 0027-8424, Document Cited by: §1.
- AMD-stability in the presence of first-order mean motion resonances. A&A 607, pp. A35. External Links: ISSN 0004-6361, Document Cited by: §4.1.
- The path to instability in compact multi-planetary systems. A&A 641, pp. A176. External Links: ISSN 0004-6361, Document Cited by: §3.2, §3.3.
- The onset of instability in resonant chains. MNRAS 494, pp. 4950–4968. External Links: ISSN 0035-8711, Document Cited by: §4.1.
- Spacing of Kepler Planets: Sculpting by Dynamical Instability. ApJ 807, pp. 44. External Links: ISSN 0004-637X, Document Cited by: §1, §3.2.
- Three-body resonance overlap in closely spaced multiple-planet systems. MNRAS 418, pp. 1043–1054. External Links: ISSN 0035-8711, Document Cited by: §3.3.
- Hybrid symplectic integrators for planetary dynamics. MNRAS 485, pp. 5490–5497. External Links: ISSN 0035-8711, Document Cited by: §2.
- High-order symplectic integrators for planetary dynamics and their implementation in REBOUND. MNRAS 489, pp. 4632–4640. External Links: ISSN 0035-8711, Document Cited by: Appendix B.
- WHFAST: a fast and unbiased implementation of a symplectic Wisdom-Holman integrator for long-term gravitational simulations. MNRAS 452, pp. 376–388. External Links: ISSN 0035-8711, Document Cited by: §2.
- Transiting Exoplanet Survey Satellite (TESS). Proc. SPIE 9143, pp. 914320. External Links: Document Cited by: §1.
- Unveiling the planet population at birth. MNRAS 503, pp. 1526–1542. External Links: ISSN 0035-8711, Document Cited by: §1, §4.4.
- Spin-Orbit Misalignment as a Driver of the Kepler Dichotomy. ApJ 830, pp. 5. External Links: ISSN 0004-637X, Document Cited by: §4.4.
- The Resilience of Kepler Systems to Stellar Obliquity. AJ 155, pp. 167. External Links: Document Cited by: §4.4.
- Evolution of protoplanetary discs with magnetically driven disc winds. A&A 596, pp. A74. External Links: ISSN 0004-6361, Document Cited by: §2, §4.3.
- A Criterion for the Onset of Chaos in Compact, Eccentric Multiplanet Systems. AJ 162, pp. 220. External Links: ISSN 0004-6256, Document Cited by: §4.1.
- First-order mean motion resonances in two-planet systems: general analysis and observed systems. MNRAS 482, pp. 530–549. External Links: ISSN 0035-8711, Document Cited by: §3.2.
- The Orbital Eccentricity of Small Planet Systems. AJ 157, pp. 61. External Links: ISSN 0004-6256, Document Cited by: §1, §4.4.
- Eccentricity from Transit Photometry: Small Planets in Kepler Multi-planet Systems Have Low Eccentricities. ApJ 808, pp. 126. External Links: ISSN 0004-637X, Document Cited by: §1.
- A Statistical Distribution Function Of Wide Applicability. Journal of Applied Mechanics 18, pp. 293–297. External Links: Document Cited by: §3.3.
- The California-Kepler Survey. V. Peas in a Pod: Planets in a Kepler Multi-planet System Are Similar in Size and Regularly Spaced. AJ 155, pp. 48. External Links: ISSN 0004-6256, Document Cited by: §1.
- The Occurrence and Architecture of Exoplanetary Systems. ARA&A 53, pp. 409. External Links: ISSN 0066-4146, Document Cited by: §1.
- Mass and Mass Scalings of Super-Earths. ApJ 874, pp. 91. External Links: ISSN 0004-637X, Document Cited by: §1.
- Planets Across Space and Time (PAST). IV. The Occurrence and Architecture of Kepler Planetary Systems as a Function of Kinematic Age Revealed by the LAMOST-Gaia-Kepler Sample. AJ 166, pp. 243. External Links: ISSN 0004-6256, Document Cited by: §5.
- How Close are Compact Multiplanet Systems to the Stability Limit?. AJ 162, pp. 55. External Links: ISSN 0004-6256, Document Cited by: §4.4.
Appendix A Intrinsic system statistics
Figures 14–17 show population statistics for the 40steep, 10straight, 20straight, and 40straight runs.
Appendix B Numerical tests